A toy manufacturer needs a piece of plastic in the shape of a right triangle with the longer leg 4 cm more than the shorter leg and the hypotenuse 8 cm more than the shorter leg. How long should the sides of the triangle​ be?
The sides are?

Found 3 solutions by jorel1380, stanbon, rfer:
Answer by jorel1380(3719)

Let n be the length of the short side. Then the longer leg is n+4, and the hypotenuse is n+8. So:
n²+(n+4)²=(n+8)²
2n²+8n+16=n²+16n+64
n²-8n-48=0
(n-12)(n+4)=0
n=12,-4
Your legs are 12, 16, and 20
☺☺☺☺

Answer by stanbon(75887)

A toy manufacturer needs a piece of plastic in the shape of a right triangle with the longer leg 4 cm more than the shorter leg and the hypotenuse 8 cm more than the shorter leg. How long should the sides of the triangle​ be?
========
shorter leg:: x
longer leg:: x + 4
hypotenuse:: x + 8
-------------------------
Equation:
x^2 + (x+4)^2 = (x+8)^2
x^2 + x^2+8x+16 = x^2+16x+64
----
x^2 - 8x - 48 = 0
Factor::
(x-12)(x+4) = 0
---
Positive solution::
x = 12 cm (shorter leg)
x+4 = 16 cm (longer leg)
x+8 = 20 cm (hypotenuse)
------------
Cheers,
Stan H.
---------------

Answer by rfer(16322)